# Optimization of Link Configuration for Satellite Communication Using Reinforcement Learning

### v1

@@ 4.2 Systemparameters
Bandwidth=SRRollout FactorSpacing $\begin{split}\text{Bandwidth}=\text{SR}\ast\text{Rollout Factor}\ast\text{Spacing Factor}\end{split}$ (4)
@@ 5.5 Training Process
Before initiating the experiments, the learning rate for the PPO was examined through a grid search. For the first experiment, the grid search included values of $31e-$ 3, $45e-$ 4, $41e-$ 4, $55e-$ 5, and $51e-$ 5. Each learning rate was evaluated using three different seeds. Across all runs, $2,000,000$ steps were executed in the transponder environment, maintaining a consistent batch size of $2,000$ for comparability. Each episode consisted of $10$ steps, serving as the termination criterion, resulting in a total of $200,000$ episodes. The best performance was obtained with a learning rate of $51e-$ 5. This learning rate was subsequently used for Experiment 1.
A similar grid search was performed for the second experiment, with learning rate values set at $31e-$ 3, $41e-$ 4, and $51e-$ 5. Each learning rate was again tested with three different seeds. For all runs, $1,000,000$ steps were carried out in the transponder environment, with a batch size of $2,000$. Episodes consisted of $100$ steps. Also for Experiment 2, a learning rate of $51e-$ 5 yielded the best performance.
@@ 6.1 Experiment 1
Figure 2 illustrates the training of the PPO model over $2,000,000$ steps, compared to the Simulated Annealing and Random Action baselines. Each PPO episode consisted of $10$ steps. To enhance visualization, all curves were smoothed. Training for the PPO model and the computation of the comparative baselines were conducted with five different seeds (00 to $4$). The figure presents the average performance across these runs, including the standard deviation for each algorithm. The maximum achievable value is $1$. It is noteworthy that only the PPO model was trained, as neither the Simulated Annealing algorithm nor the Random Action model can be trained. These baselines were plotted to provide a benchmark for assessing the PPO’s performance.
@@ 6.2 Experiment 2
Experiment $2$ differs from Experiment $1$ in that it uses Action Space $2$, defined such that a parameter value can be modified from a link in a single step. The parameters available for modification include center frequency, EIRP, and the MOD-FEC combination. As in Experiment 1, PPO was compared with the Simulated Annealing and Random Action baselines. The selected learning rate for PPO was $51e-$ 5, as this yielded the best performance during the grid search for Action Space $2$.

### v2

@@ 4.2 Systemparameters
$\begin{split}\text{Bandwidth}=\text{SR}\ast\text{Rollout Factor}\ast\text{Spacing Factor}\end{split}$ (4)
@@ 5.5 Training Process
Before initiating the experiments, the learning rate for the PPO was examined through a grid search. For the first experiment, the grid search included values of 1e31e-3, 5e45e-4, 1e41e-4, 5e55e-5, and 1e51e-5. Each learning rate was evaluated using three different seeds. Across all runs, $2,000,000$ steps were executed in the transponder environment, maintaining a consistent batch size of $2,000$ for comparability. Each episode consisted of $10$ steps, serving as the termination criterion, resulting in a total of $200,000$ episodes. The best performance was obtained with a learning rate of 1e51e-5. This learning rate was subsequently used for Experiment 1.
A similar grid search was performed for the second experiment, with learning rate values set at 1e31e-3, 1e41e-4, and 1e51e-5. Each learning rate was again tested with three different seeds. For all runs, $1,000,000$ steps were carried out in the transponder environment, with a batch size of $2,000$. Episodes consisted of $100$ steps. Also for Experiment 2, a learning rate of 1e51e-5 yielded the best performance.
@@ 6.1 Experiment 1
Figure 2 illustrates the training of the PPO model over $2,000,000$ steps, compared to the Simulated Annealing and Random Action baselines. Each PPO episode consisted of $10$ steps. To enhance visualization, all curves were smoothed. Training for the PPO model and the computation of the comparative baselines were conducted with five different seeds (0 to $4$). The figure presents the average performance across these runs, including the standard deviation for each algorithm. The maximum achievable value is $1$. It is noteworthy that only the PPO model was trained, as neither the Simulated Annealing algorithm nor the Random Action model can be trained. These baselines were plotted to provide a benchmark for assessing the PPO’s performance.
@@ 6.2 Experiment 2
Experiment $2$ differs from Experiment $1$ in that it uses Action Space $2$, defined such that a parameter value can be modified from a link in a single step. The parameters available for modification include center frequency, EIRP, and the MOD-FEC combination. As in Experiment 1, PPO was compared with the Simulated Annealing and Random Action baselines. The selected learning rate for PPO was 1e51e-5, as this yielded the best performance during the grid search for Action Space $2$.